# Learning to summarize user information for personalized reinforcement learning from human feedback

### v2

@@ Abstract
As everyday use cases of large language model (LLM) AI assistants have expanded, it is becoming increasingly important to personalize responses to align to different users’ preferences and goals. While reinforcement learning from human feedback (RLHF) is effective at improving LLMs to be generally more helpful and fluent, it does not account for variability across users, as it models the entire user population with a single reward model, meaning it assumes that everyone’s preferences are the same. We present a novel framework, Preference Learning Using Summarization (PLUS), that uses reinforcement learning (RL) to learn to produce text-based summaries of each user’s preferences, characteristics, and past conversations. These summaries condition the reward model, enabling it to make personalized predictions about the types of responses valued by each user. Both the user-summarization model and reward model are trained simultaneously, creating an online co-adaptation loop. We show that in contrast to the standard Bradley–Terry model, summaries produced by PLUS capture diverse aspects of user preferences, achieving a 11–77% improvement in reward model accuracy. Key strengths of PLUS are: (1) robust performance with new users and conversation topics, achieving a 25% improvement over the best personalized RLHF technique; (2) zero-shot personalization with state-of-the-art proprietary models like GPT-4 (e.g., PLUS-summary-conditioned responses achieved a 72% win rate compared to 28% for default GPT-4o); (3) learning from flexible user contexts beyond preference labels, and (4) interpretable representation of users, enabling greater transparency and user control in pluralistic LLM alignment.
@@ 1 Introduction
Large language models (LLMs) have become a daily part of life for over a billion people, with Gemini reaching at least 450 million monthly users by July [13] and ChatGPT’s daily users approaching a billion [6]. Users interact with LLMs in various ways, from asking for domain-specific help to engaging in open-ended storytelling and content creation[34]. Reinforcement Learning from Human Feedback (RLHF  [17]) is a critical component of modern LLMs deployed in production systems, because it ensures that they are not only functional (i.e., usefully follow user instructions) but also aligned with constitutional values such as safety and helpfulness [1]. The standard Bradley-Terry-Luce (BTL) reward model [2] used in existing RLHF algorithms assumes all users share the same preferences over how the LLM should respond. This technique is incapable of creating systems that can adapt to users’ unique needs and preferences, as a single reward model is used to represent all users.
@@
Figure 1: While standard RLHF techniques fail to capture user variability, PLUS trains both a summarizer and reward model in an online co-adaptive framework to learn summaries $z$ useful for predicting diverse preferences. Italicized texts are actual outputs by GPT-4o and PLUS showing the effects of summaries on personalization.
@@
However, real users have diverse, potentially conflicting, preferences  [12], and current methods are unable to accommodate these differences. In fact, when the standard BTL model is applied to data with conflicting preferences, it will either ignore the preferences of minority groups [23], or learn a reward model that is inaccurate for all groups of users [18]. This has led to recent calls to develop techniques for pluralistic alignment [26, 14, 23], to enable LLMs to be personalized to heterogeneous user preferences. Given the widespread use of LLMs and the diverse populations of users they serve, being able to model and adhere to more than one set of preferences is critical for making effective, accurate assistants that are capable of personalizing to each individual’s unique needs. One way to achieve pluralistic alignment is by developing a reward model conditioned on user variability, rather than relying on a single BTL model [23, 32, 18] (e.g.,  [18] learns a latent user embedding from the past preference labels). However, compressing rich text into a vector can lead to performance loss [4]. As yet, we are missing a highly performant, public RLHF technique capable of modeling the preferences of diverse users.
@@
A more natural approach which retains the strengths of LLMs to reason with text, is to learn text-based user summaries to act as the latent user variable. We propose Preference Learning Using Summaries (PLUS) as a general RLHF framework that learns to summarize important information about the user using reinforcement learning (RL) fine-tuning and conditions the reward model on these user summaries. PLUS consists of two co-dependent training phases: (1) a summarizer trained with Proximal Policy Optimization (PPO) [20], and (2) a summary-conditioned reward model, which provides the reward signal for training the summarizer and makes predictions about user preferences.
@@
We carefully compare our approach to several alternative implementations. First, we check whether we can simply prompt an LLM to generate user summaries without fine-tuning the summarizer. Our experiments show that automatic summaries (even from state-of-the-art GPT-4) often miss key details necessary to capture user preferences. Another alternative is to condition the reward model on the user’s entire conversation history, without summarizing it. However, both approaches fall short compared to concise, summary-conditioned reward models when conversation topics change. An additional benefit of concise summaries compared to existing techniques is that the summaries are easier to review and edit, improving transparency and trust in LLM assistants. We show that such benefits extend to complex real-world datasets, like PRISM [12], where diverse user preferences can be expressed through flexible textual summaries and used to enable zero-shot personalization. In summary, our main contributions are:
@@
2. Empirical advantages of PLUS on benchmarks from prior state-of-the-art work [18]: Pets and UltraFeedback [5]. PLUS improves reward model accuracy by 11-77% compared to BTL, is robust to topic shifts, and can handle unstructured user contexts.
3. Scaling to a dataset in the wild. We are one of the first to report reward modeling results with PRISM [12], a pluralistic dataset of 1,500 users across 75 countries and 20 LLMs.
4. Enabling zero-shot personalization via PLUS-generated summaries. PLUS can generalize to unseen users of the PRISM dataset and the learned summaries can be used to personalize responses of strong proprietary models without additional finetuning (PLUS summary-conditioned responses achieve a 72% win rate over the default GPT-4o responses).
@@ 2 Related Work
Reinforcement Learning from Human Feedback (RLHF). Our work builds on reinforcement learning from human feedback (RLHF)  [17], which models binary human preferences using the Bradley-Terry-Luce (BTL) model [2]. The BTL model assumes the user provides preference ratings in terms of which of two responses they like better; it then learns a reward model (RM) that estimates the magnitude of the user’s preferences using this data. This is then used to train an LLM policy to optimize responses based on the reward signal. RLHF has been effective in refining language models’ responses to better align with user-preferred attributes, like helpfulness and truthfulness [1]. However, existing RLHF works [35, 1, 17, 23, 19] have assumed a single reward model to capture all user data, which may be misaligned with heterogeneous users.
@@
Personalized & Pluralistic RLHF. Increasingly, more works [25, 11, 12, 32, 18, 14, 3, 15] have pointed out the limitations of existing BTL reward models in accommodating diverse underlying preferences, which may conflict with each other in terms of what individual users consider ideal responses. These differences may surface along various dimensions, from writing style to value alignment. Despite algorithmic advances in pluralistic alignment, large-scale human datasets are still limited. PRISM [12] makes an effort to provide a heterogeneous user dataset consisting of 1,500 participants from 75 countries interacting with 21 different LLMs (contributing around 9,000 unique conversations). However, so far we are not aware of work that has successfully published reward modeling results on PRISM, pointing to the difficulty of modeling diverse and heterogeneous preferences with sparse data.
@@
Broadly, works on pluralistic RLHF have focused on either of two main aspects: (1) personalizing reward models [23, 18, 32, 15, 3] and (2) personalizing LLM responses with a user model [14]. Reward (user) modeling serves as the backbone of many existing algorithms for LLM alignment, since reward model accuracy has direct implications for downstream tasks, such as RLHF finetuning and best-of-N sampling [22, 8, 16]. Notably,  Poddar et al. [18] introduces Variational Preference Learning (VPL) which uses a variational autoencoder to learn user-specific embeddings and conditions the reward model on these latent variables.  Siththaranjan et al. [23], Zhao et al. [32], Chen et al. [3] provide algorithms for accommodating diverse user groups through methods such as Distributional Preference Learning (DPL), dispersion index, and shared LoRA adaptation. However, the key difference between prior works and ours is that these methods represent users as embedding vectors or (linear) weights, which reduces the power and interpretability of the underlying personalization mechanism. In contrast, we leverage the text-generation capabilities of LLMs to represent diverse user preferences through summaries.
@@
LLM finetuning for summarization. Finetuning LLMs for improved and domain-specific summarization has long been of interest [27, 35, 31, 28, 24, 29]. Most closely related to our work is by  Wu et al. [29] which demonstrate that learning user summaries can be effective for personalizing recommendation systems, when you assume access to a fixed user classifier. In contrast, we focus on the online co-adaptive loop of training both the summarizer and reward model, which serves as the key for enabling the reward model to make use of the learned summaries for capturing nuanced user preferences.
@@ 3 Preliminaries
Bradley-Terry-Luce reward model. Our work focuses on the reward modeling stage of the standard RLHF pipeline [17]. Given a dataset of annotated preferences $\mathcal{D}=\{(s_{A}^{i},s_{B}^{i},\mathbbm{1}\{s_{A}^{i}\succ s_{B}^{i}\})\}_{i=1}^{N}$, where $s_{A}=[x,y_{A}]$ and $y_{B}=[x,y_{B}]$ are two responses $y_{A},y_{B}$ over the same prompt $x$, the standard Bradley-Terry-Luce (BTL) model [2] assumes an underlying reward function $r(s)$. This reward function determines the likelihood of the preference label $\mathbbm{1}\{s_{A}\succ s_{B}\}$ as:
@@
User-conditioned reward model. Existing work in personalized RLHF [14, 18] has demonstrated the effectiveness of extending the standard BTL reward model into a user-conditioned reward model:
@@
However, in most real-world scenarios, datasets do not contain user identifiers $z$. This is partly because such identifiers may be unavailable due to privacy issue, and partly because conditioning only on users observed in the dataset makes it difficult to generalize to new users. Therefore, we can instead use context $c$, such as past conversations, to represent the current user in each data point111 Poddar et al. [18] requires that the context consist of a user’s past chosen and rejected responses to different prompts $c=\{(s_{A}^{j},s_{B}^{j})\}_{j=1}^{m}$. In contrast, PLUS can incorporate more flexible inputs like past conversations without explicit preference labels and system strings.. This leads to a dataset of user-conditioned annotated preferences $\mathcal{D}^{\prime}=\{(s_{A}^{i},s_{B}^{i},c^{i},z^{i},\mathbbm{1}\{s_{A}^{i}\succ s_{B}^{i}|z^{i}\})\}_{i=1}^{N}$ where $z^{i}$ are unknown. For notational convenience, we assume $s_{A}^{i}$ is always the preferred one by user $z^{i}$, and the dataset can be simplified to $\mathcal{D}^{\prime}=\{(s_{A}^{i},s_{B}^{i},c^{i})\}_{i=1}^{N}$.
@@ 4 Preference Learning Using Summarization: PLUS
While prior works [14, 18] formulate the latent user representation $z$ as an embedding vector, we propose learning $z$ in natural language as a more natural way to leverage the advanced text-based reasoning and generation capabilities of LLMs. To be more specific, we propose to use a summarizer $\pi_{\theta}$ to generate a concise representation in natural language based on the context $c$ of each user, $z\sim\pi_{\theta}(\cdot|c)$. Then, the parametrized reward model $r_{\phi}(\cdot|z)$ will compute the reward conditioned on this learned summary $z$.
@@
To handle the co-dependency in Eq. 3, we design a two-stage sequential (batch) training process that alternates between training one model while keeping the other fixed. By allowing the summarizer and the reward model to co-adapt, the summarizer learns to extract information that is most useful for improving the predictive accuracy of the reward model, while the reward model is trained on updated summaries to more accurately capture user preferences. This process is computationally efficient, since the generated summaries always target the most recent iteration of the reward model.
Training the reward model. Once the summarizer is fixed, the reward model is trained to minimize the negative log-likelihood loss (Eq. 3) using pairs of chosen and rejected responses, conditioned on user summaries sampled from the summarizer policy.
Training the summarizer. Unlike the reward model, a summarizer’s optimal weights cannot be learned directly through supervised learning, because whether a summary is effective for predicting preferences can only be assessed once the full summary has been sampled. Instead, RL can optimize the summarizer at the token level by assigning credit for the full-trajectory return across long token sequences $(z_{1},z_{2},...,z_{t\leq H})$, as demonstrated by prior successes in using RL to improve summarization [27, 35]. Thus, we formulate generating the summary into an RL post-training problem, with the objective in Eq. 4. We start with a default LLM as a reference policy $\pi_{\theta}$, sample summaries (trajectories) from the current $\pi_{\theta}$, estimate their returns (as negative losses of the reward model), and update the weights $\theta$ in a direction that maximizes estimated returns.
@@
PPO training for handling non-stationarity. As the reward model for predicting user preferences is continuously updated, the RL rewards for training the summarizer become non-stationary, as in a multi-agent RL (MARL) problem. Therefore, while different RL algorithms could be considered for learning an effective summarizer, we choose PPO [20], as it is a state-of-the-art MARL algorithm and performs well even under non-stationary rewards [7, 30]. In addition to the summary-level rewards $R=\sum_{t=0}^{H}\gamma^{t}R_{t}$, we calculate token-level advantages using the generalized advantage estimator (GAE) [21]:
@@
We finetuned different reward models (Qwen2.5-0.5B-Instruct, Qwen2.5-3B-Instruct, and Llama3.2-3B-Instruct) and summarizers (Qwen2.5-3B-Instruct or Llama3.2-3B-Instruct). We pair each reward model with a summarizer from the same model family. The models are fine-tuned using low-rank adaptation (LoRA) [9] with distributed computation across multiple nodes (we used 4 x H200 GPUs). The co-adaptation pipeline in Alg. 1 is built on top of PPO training framework from OpenRLHF [10]. We follow the baseline implementations for BTL, VPL, and DPL by  Poddar et al. [18]. More details are in Appendix H.

### v3

@@ Abstract
As everyday use cases of large language model (LLM) AI assistants have expanded, it is becoming increasingly important to personalize responses to align to different users’ preferences and goals. While reinforcement learning from human feedback (RLHF) is effective at improving LLMs to be generally more helpful and fluent, it does not account for variability across users, as it models the entire user population with a single reward model, meaning it assumes that everyone’s preferences are the same. We present a novel framework, Preference Learning Using Summarization (PLUS), that uses reinforcement learning (RL) to learn to produce text-based summaries of each user’s preferences, characteristics, and past conversations. These summaries condition the reward model, enabling it to make personalized predictions about the types of responses valued by each user. Both the user-summarization model and reward model are trained simultaneously, creating an online co-adaptation loop. We show that in contrast to the standard Bradley–Terry model, summaries produced by PLUS capture diverse aspects of user preferences, achieving a 11–77% improvement in reward model accuracy. Key strengths of PLUS are: (1) robust performance with new users and conversation topics, achieving a 25% improvement over the best personalized reward model technique used for RLHF; (2) personalization with strong proprietary models like GPT-4 without further fine-tuning (e.g., PLUS-summary-conditioned responses achieved a 72% win rate compared to 28% for default GPT-4o); (3) learning from flexible user contexts beyond preference labels, and (4) interpretable representation of users, enabling greater transparency and user control in pluralistic LLM alignment.
@@ 1 Introduction
Large language models (LLMs) have become a daily part of life for over a billion people, with Gemini reaching at least 450 million monthly users by July (Lee, 2025) and ChatGPT’s daily users approaching a billion (Curry, 2025). Users interact with LLMs in various ways, from asking for domain-specific help to engaging in open-ended storytelling and content creation(Zheng et al., 2024). Reinforcement Learning from Human Feedback (RLHF  (Ouyang et al., 2022)) is a critical component of modern LLMs deployed in production systems, because it ensures that they are not only functional (i.e., usefully follow user instructions) but also aligned with constitutional values such as safety and helpfulness (Bai et al., 2022). The standard Bradley-Terry-Luce (BTL) reward model (Bradley and Terry, 1952) used in existing RLHF algorithms learns a reward function based on majority preferences: If two users have opposite preferences about the response to a particular prompt, and no further context is provided, the BTL model cannot learn to disambiguate between different users’ preferences.
@@
Figure 1: Caution: This content may reflect particular beliefs. While standard RLHF techniques fail to capture user variability, PLUS trains both a summarizer and reward model in an online co-adaptive framework to learn summaries $z$ useful for predicting diverse preferences. Italicized texts are actual outputs by GPT-4o and PLUS showing the effects of summaries on personalization.
@@
However, real users have diverse, potentially conflicting, preferences  (Kirk et al., 2024), and current methods are unable to accommodate these differences. In fact, when the standard BTL model is applied to data with conflicting preferences, it will either ignore the preferences of minority groups (Siththaranjan et al., 2024), or learn a reward model that is inaccurate for all groups of users (Poddar et al., 2024). This has led to recent calls to develop techniques for pluralistic alignment Sorensen et al. (2024); Li et al. (2024); Siththaranjan et al. (2024), to enable LLMs to be personalized to heterogeneous user preferences. Given the widespread use of LLMs and the diverse populations of users they serve, being able to model and adhere to more than one set of preferences is critical for making effective, accurate assistants that are capable of personalizing to each individual’s unique needs. One way to achieve pluralistic alignment is by developing a reward model conditioned on user variability, rather than relying on a single BTL model (Siththaranjan et al., 2024; Zhao et al., 2024; Poddar et al., 2024) (e.g.,  (Poddar et al., 2024) learns a latent user embedding from the past preference labels). However, compressing rich text into a vector can lead to performance loss (Cideron et al., 2022). As yet, we are missing a highly performant, public RLHF technique capable of modeling the preferences of diverse users.
@@
A more natural approach which retains the strengths of LLMs to reason with text, is to learn text-based user summaries to act as the latent user variable. We propose Preference Learning Using Summaries (PLUS) as a general RLHF framework that learns to summarize important information about the user using reinforcement learning (RL) fine-tuning and conditions the reward model on these user summaries. PLUS consists of two co-dependent training phases: (1) a summarizer trained with Proximal Policy Optimization (PPO) (Schulman et al., 2017), and (2) a summary-conditioned reward model, which provides the reward signal for training the summarizer and makes predictions about user preferences.
@@
We carefully compare our approach to several alternative implementations. First, we check whether we can simply prompt an LLM to generate user summaries without fine-tuning the summarizer. Our experiments show that automatic summaries (even from GPT-4) often miss key details necessary to capture user preferences. Another alternative is to condition the reward model on the user’s entire conversation history, without summarizing it. However, both approaches fall short compared to concise, summary-conditioned reward models when conversation topics change. An additional benefit of concise summaries compared to existing techniques is that the summaries are easier to review and edit, improving transparency and trust in LLM assistants. We show that such benefits extend to complex real-world datasets, like PRISM (Kirk et al., 2024), where diverse user preferences can be expressed through flexible textual summaries and used to enable personalization of strong, proprietary models without additional fine-tuning. In summary, our main contributions are:
@@
2. Empirical advantages of PLUS on benchmarks from prior state-of-the-art work (Poddar et al., 2024): Pets and UltraFeedback (Cui et al., 2024). PLUS improves reward model accuracy by 11-77% compared to BTL, is robust to topic shifts, and can handle unstructured user contexts.
3. Scaling to a dataset in the wild. We are one of the first to report reward modeling results with PRISM (Kirk et al., 2024), a pluralistic dataset of 1,500 users across 75 countries and 20 LLMs.
4. Enabling personalization of strong, proprietary models via PLUS-generated summaries. PLUS can generalize to unseen users of the PRISM dataset and the learned summaries can be used to personalize responses of strong proprietary models without additional finetuning (PLUS summary-conditioned responses achieve a 72% win rate over the default GPT-4o responses).
@@ 2 Related Work
Reinforcement Learning from Human Feedback (RLHF). Our work builds on reinforcement learning from human feedback (RLHF)  (Ouyang et al., 2022), which models binary human preferences using the Bradley-Terry-Luce (BTL) model (Bradley and Terry, 1952). The BTL model assumes the user provides preference ratings in terms of which of two responses they like better; it then learns a reward model (RM) that estimates the magnitude of the user’s preferences using this data. This is then used to train an LLM policy to optimize responses based on the reward signal. RLHF has been effective in refining language models’ responses to better align with user-preferred attributes, like helpfulness and truthfulness (Bai et al., 2022). However, existing RLHF works (Ziegler et al., 2020; Bai et al., 2022; Ouyang et al., 2022; Siththaranjan et al., 2024; Rafailov et al., 2024) have assumed a single reward model to capture all user data, which may be misaligned with heterogeneous users.
@@
Personalized & Pluralistic RLHF. Increasingly, more works (Sorensen et al., 2023; Kirk et al., 2023; 2024; Zhao et al., 2024; Poddar et al., 2024; Li et al., 2024; Chen et al., 2025; Liu et al., 2025) have pointed out the limitations of existing BTL reward models in accommodating diverse underlying preferences, which may conflict with each other in terms of what individual users consider ideal responses. These differences may surface along various dimensions, from writing style to value alignment. Despite algorithmic advances in pluralistic alignment, large-scale human datasets are still limited. PRISM (Kirk et al., 2024) makes an effort to provide a heterogeneous user dataset consisting of 1,500 participants from 75 countries interacting with 21 different LLMs (contributing around 9,000 unique conversations). However, so far we are not aware of work that has successfully published reward modeling results on PRISM, pointing to the difficulty of modeling diverse and heterogeneous preferences with sparse data.
@@
Broadly, works on pluralistic RLHF have focused on either of two main aspects: (1) personalizing reward models (Siththaranjan et al., 2024; Poddar et al., 2024; Zhao et al., 2024; Liu et al., 2025; Chen et al., 2025) and (2) personalizing LLM responses with a user model (Li et al., 2024). Reward (user) modeling serves as the backbone of many existing algorithms for LLM alignment, since reward model accuracy has direct implications for downstream tasks, such as RLHF finetuning and best-of-N sampling (Shen et al., 2023; Gao et al., 2022; Meng et al., 2024). Notably,  Poddar et al. (2024) introduces Variational Preference Learning (VPL) which uses a variational autoencoder to learn user-specific embeddings and conditions the reward model on these latent variables.  Siththaranjan et al. (2024); Zhao et al. (2024); Chen et al. (2025) provide algorithms for accommodating diverse user groups through methods such as Distributional Preference Learning (DPL), dispersion index, and shared LoRA adaptation. However, the key difference between prior works and ours is that these methods represent users as embedding vectors or (linear) weights, which reduces the power and interpretability of the underlying personalization mechanism. In contrast, we leverage the text-generation capabilities of LLMs to represent diverse user preferences through summaries.
@@
LLM finetuning for summarization. Finetuning LLMs for improved and domain-specific summarization has long been of interest (Stiennon et al., 2022; Ziegler et al., 2020; Zhang et al., 2023; Van Veen et al., 2023; Song et al., 2025; Wu et al., 2025). Most closely related to our work is by  Wu et al. (2025) which demonstrate that learning user summaries can be effective for personalizing recommendation systems, when you assume access to a fixed user classifier. In contrast, we focus on the online co-adaptive loop of training both the summarizer and reward model, which serves as the key for enabling the reward model to make use of the learned summaries for capturing nuanced user preferences.
@@ 3 Preliminaries
Bradley-Terry-Luce reward model. Our work focuses on the reward modeling stage of the standard RLHF pipeline (Ouyang et al., 2022). Given a dataset of annotated preferences $\mathcal{D}=\{(s_{A}^{i},s_{B}^{i},\mathbbm{1}\{s_{A}^{i}\succ s_{B}^{i}\})\}_{i=1}^{N}$, where $s_{A}=[x,y_{A}]$ and $y_{B}=[x,y_{B}]$ are two responses $y_{A},y_{B}$ over the same prompt $x$, the standard Bradley-Terry-Luce (BTL) model (Bradley and Terry, 1952) assumes an underlying reward function $r(s)$. This reward function determines the likelihood of the preference label $\mathbbm{1}\{s_{A}\succ s_{B}\}$ as:
@@
User-conditioned reward model. Existing work in personalized RLHF (Li et al., 2024; Poddar et al., 2024) has demonstrated the effectiveness of extending the standard BTL reward model into a user-conditioned reward model:
@@
However, in most real-world scenarios, datasets do not contain user identifiers $z$. This is partly because such identifiers may be unavailable due to privacy issue, and partly because conditioning only on users observed in the dataset makes it difficult to generalize to new users. Therefore, we can instead use context $c$, such as past conversations, to represent the current user in each data point111 Poddar et al. (2024) requires that the context consist of a user’s past chosen and rejected responses to different prompts $c=\{(s_{A}^{j},s_{B}^{j})\}_{j=1}^{m}$. In contrast, PLUS can incorporate more flexible inputs like past conversations without explicit preference labels and system strings.. This leads to a dataset of user-conditioned annotated preferences $\mathcal{D}^{\prime}=\{(s_{A}^{i},s_{B}^{i},c^{i},z^{i},\mathbbm{1}\{s_{A}^{i}\succ s_{B}^{i}|z^{i}\})\}_{i=1}^{N}$ where $z^{i}$ are unknown. For notational convenience, we assume $s_{A}^{i}$ is always the preferred one by user $z^{i}$, and the dataset can be simplified to $\mathcal{D}^{\prime}=\{(s_{A}^{i},s_{B}^{i},c^{i})\}_{i=1}^{N}$.
@@ 4 Preference Learning Using Summarization: PLUS
While prior works (Li et al., 2024; Poddar et al., 2024) formulate the latent user representation $z$ as an embedding vector, we propose learning $z$ in natural language as a more natural way to leverage the advanced text-based reasoning and generation capabilities of LLMs. To be more specific, we propose to use a summarizer $\pi_{\theta}$ to generate a concise representation in natural language based on the context $c$ of each user, $z\sim\pi_{\theta}(\cdot|c)$. Then, the parametrized reward model $r_{\phi}(\cdot|z)$ will compute the reward conditioned on this learned summary $z$.
@@
To handle the co-dependency in Eq. 3, we design a training process that continues to alternate between training one model while keeping the other fixed. By allowing the summarizer and the reward model to co-adapt, the summarizer learns to extract information that is most useful for improving the predictive accuracy of the reward model, while the reward model is trained on updated summaries to more accurately capture user preferences. This process is computationally efficient, since the generated summaries always target the most recent iteration of the reward model.
Training the reward model. When training the reward model, the summarizer weights are fixed, and the reward model is trained to minimize the negative log-likelihood loss (Eq. 3) using pairs of chosen and rejected responses, conditioned on user summaries sampled from the summarizer policy. Then, we repeat the step of training the summarizer using the newly obtained reward model weights.
@@
Training the summarizer. Unlike the reward model, a summarizer’s optimal weights cannot be learned directly through supervised learning, because whether a summary is effective for predicting preferences can only be assessed once the full summary has been sampled. Instead, RL can optimize the summarizer at the token level by assigning credit for the full-trajectory return across long token sequences $(z_{1},z_{2},...,z_{t\leq H})$, as demonstrated by prior successes in using RL to improve summarization (Stiennon et al., 2022; Ziegler et al., 2020). Thus, we formulate generating the summary into an RL post-training problem, with the objective in Eq. 4. We start with a default LLM as a reference policy $\pi_{\theta}$, sample summaries (trajectories) from the current $\pi_{\theta}$, estimate their returns (as negative losses of the reward model), and update the weights $\theta$ in a direction that maximizes estimated returns.
@@
PPO training for handling non-stationarity. As the reward model for predicting user preferences is continuously updated, the RL rewards for training the summarizer become non-stationary, as in a multi-agent RL (MARL) problem. Therefore, while different RL algorithms could be considered for learning an effective summarizer, we choose PPO (Schulman et al., 2017), as it is a state-of-the-art MARL algorithm and performs well even under non-stationary rewards (de Witt et al., 2020; Yu et al., 2022). In addition to the summary-level rewards $R=\sum_{t=0}^{H}\gamma^{t}R_{t}$, we calculate token-level advantages using the generalized advantage estimator (GAE) (Schulman et al., 2018):
@@
Why learning to summarize is important.
An untrained summarizer can lead to summaries that focus on past content rather than the general style of conversations the user prefers to have. Therefore, the obtained summaries may not provide predictive signal for the reward model. Alternatively, one could include all prior preference labels from the user as in-context learning (ICL) examples, and train the reward model conditioned on such examples (i.e., the full list of examples acting as the conditioning variable $z$). However, the large input space of ICL may make reward learning more difficult. This shows up in our empirical results, where ICL underperforms, especially when trying to generalize to new conversation topics. While in theory ICL can enable personalization, it may be sample-inefficient and computationally expensive to train and use for inference due to the increased context length.
@@
We finetuned different reward models (Qwen2.5-0.5B-Instruct, Qwen2.5-3B-Instruct, and Llama3.2-3B-Instruct) and summarizers (Qwen2.5-3B-Instruct or Llama3.2-3B-Instruct). We paired each reward model with a summarizer from the same model family. We used 2-4 x H200 GPUs for model finetuning. The co-adaptation pipeline in Alg. 1 is built on top of PPO training framework from OpenRLHF (Hu et al., 2024), and we also used the reward model training code from OpenRLHF. For other baselines, namely VPL, DPL, and BTL (for Pets), we used the implementations by  Poddar et al. (2024). More details are in Appendix I.